OpenStudy (anonymous):

Part 1: use the zero product property to set up two equations that will lead to solutions to the original equation. Part 2: Use a reciprocal identity to express the equation involving secant in terms of sine, cosine, or tangent. Part 3: Solve each of the two equations in Part 1 for x, giving all solutions to the equation. equation below

OpenStudy (anonymous):

\[(\cos x -\sqrt{2}/2)(\sec x -1)=0\]

OpenStudy (nory):

By the zero product property: cos x - √2/2 = 0 and: sec x - 1 = 0.